Dome shutter

This page provides updates on the progress of the recovery efforts following the dome shutter incident on January 5 2013.

Friday 01 February 01 2013

Observations restarted today. Thanks to everyone involved. The shutter is still being operated carefully and tests will continue to find the causes of the problem.

Tuesday 29 January 29 2013

The team was able to close the shutter on Sunday. Congratulations to everyone involved. Because of the weather and the still precarious state of the shutter, the exchange to WIRCamis canceled. The plan is now to exchange to MegaCam tomorrow. Plans are also being developed to enable a safe opening of the shutter.

Friday January 25 2013

Power was brought back up yesterday and is holding up. Instruments are being checked and getting ready for observing. The computer system is partially back online. Right now, the plan is to try to close the dome shutter on Sunday and do an exchange to WIRCam

Wednesday January 23 2013

Preparations for the power up continue at the summit and in Waimea. Observations are still suspended.

Monday January 21 2013

The cleanup and drainage on all the floors was completed today. Power is still out at the summit but the efforts to bring it back continued today. Several breakers were cleaned and dried. Observations are still halted.

Sunday January 20 2013

There has been a major water leak on the 4th floor caused by a frozen and rupptured water line going to the humidifier room. Standing water was found at least 1 inch deep on the 4th floor, with considerable leakage to all floors below. Building power has been significantly disrupted. Work is underway to repair and insulate the water line and to limit power. All instruments have been shutdown. Observations have been halted until further notice.

Friday January 18 2013

The team appears to have identified problems associated with scarping contact on the bottom of the shutter rail near or below the spring line on the side of the dome opposite to the shutter.

Thursday January 17 2013

After several attempts, the shutter was moved about a foot in the closed direction. This came after a long day of work at the summit. Although the shutter moved, we still have no good understanding of the mecanism that got the shutter to get stuck.

Wednesday January 16 2013

Troubleshooting of the problem continues. The engineers plan to use a fiber scope under the skin to to look at the state of the side rollers. The fiber scope comes equipped with a camera so images can be saved for later review.

Monday January 14 2013

Observations resumed Sunday. Thanks for everyone involved in this effort. Observations will continue as long as the weather looks good. Plans for further work this week were developed and a schedule was build to coordinate troubleshooting.

Saturday January 12 2013

More troubleshooting is being done over the week-end. With good weather coming, it is possible that the telescope be de-winterized tomorrow for observations to resume Sunday evening. If observations resume, they will be on a temporary basis and will be stopped at the sign of inclement weather.

Friday January 11 2013

More troubleshooting was done on the shutter. Two short moves, one open and one close, were commanded and video were gathered of each of the pinions and motors. This test revealed a possible cause for the problem. This will be investigated next week.

Wednesday January 8 2013

With bad weather forcasted, the telescope was winterized and MegaCam was shutdown. Observations are suspended until further notice. Some troubleshooting was done to see if the problem was related to the shutter brakes. This didn't turn out to be the case.

Saturday January 5 2012

The shutter got stuck open at 23:54:44. A routine close due to high humidity brought the dome shutter to a stop in an almost fully open position. Work on this problem started immediately and is highest priority. Observations will continue until weather becomes a concern.